# Unifying Decoherence and Phase Evolution in Mixed Quantum–Classical Dynamics through Exact Factorization

### v1

@@ Abstract
We propose new mixed quantum–classical equations of motion within the exact factorization framework to describe electronic coherence and phase evolution simultaneously. The derivation reveals that not only a projected quantum momentum correction, recently identified by Arribas and Maitra [Phys. Rev. Lett. 133, 233201 (2024)], but also a phase-correction appears within the order of $\hbar$, providing a unified and rigorous account of electronic coherence and phase evolution as well as their effect on the nuclear force. Benchmark tests on one- and two-dimensional model systems demonstrate that the new formulations capture key nonadiabatic features with high accuracy.
@@
Understanding the coupled motion of electrons and nuclei in molecules remains one of the central challenges in theoretical chemistry and chemical physics. Fully quantum treatments of electron–nuclear dynamics are computationally prohibitive for systems of realistic size, owing to the scaling of the many-body molecular wavefunction. To overcome this limitation, mixed quantum–classical (MQC) methods [1, 2, 3] where the electrons are treated quantum mechanically while nuclei follow classical trajectories have become indispensable tools for simulating photochemical processes, charge and energy transfer, and ultrafast dynamics in molecules and materials [4, 5, 6, 7, 8, 9]. Among MQC methods, Ehrenfest dynamics [1] and trajectory-based surface hopping [2] are the most widely applied. Despite their utility, they suffer from two fundamental shortcomings: the absence of a reliable mechanism for electronic decoherence, and the lack of a rigorous treatment of phase evolution between adiabatic states. Numerous correction schemes have been proposed, often guided by semiclassical arguments, empirical adjustments, intuition from wave-packet models, or mapping of variables [10, 11, 12, 13, 14, 15, 16, 17].
@@
The exact factorization (XF) formalism provides a rigorous foundation for electron–nuclear dynamics by expressing the molecular wavefunction as a product of a nuclear wavefunction and a conditional electronic state [18, 19]. This leads to coupled electronic and nuclear equations of motion that, in principle, include the full electron–nuclear correlation (ENC) effects. XF-based approaches have therefore attracted attention as a systematic framework for deriving improved MQC dynamics [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. In recent work, Arribas and Maitra showed that a projected quantum momentum (PQM) correction naturally arises from the XF formalism and plays a central role in describing electronic coherence [32].
@@
In this work, we derive new mixed quantum–classical equations of motion from the exact factorization formalism, revealing an additional phase-correction term that accompanies the PQM correction of comparable or smaller order in $\hbar$. This previously unidentified term governs the proper phase evolution of the Born–Oppenheimer (BO) coefficients. The combined inclusion of PQM and phase corrections provides a balanced and rigorous description of both decoherence and phase dynamics without relying on semiclassical approximations or heuristic arguments, offering deeper conceptual insight into ENC and a practical route to accurate molecular quantum dynamics. We validate the equations through benchmark calculations for the double-arch geometry (DAG) [14] and two-dimensional nonseparable (2DNS) [15] models, which demonstrate that both corrections are essential for reproducing characteristic nonadiabatic features such as Stückelberg oscillations, intermediate-time electronic coherence, and the correct nuclear-density distributions.
@@
where $\hat{H}_{\mathrm{BO}}$ is the BO Hamiltonian, and $\textbf{A}_{\nu}({\underline{\underline{\textbf{R}}}},t)=\langle\Phi_{\underline{\underline{\textbf{R}}}}|-i\hbar\nabla_{\nu}\Phi_{\underline{\underline{\textbf{R}}}}\rangle_{e}$ and $\tilde{\epsilon}({\underline{\underline{\textbf{R}}}},t)=\langle\Phi_{\underline{\underline{\textbf{R}}}}(t)|(\hat{H}_{\mathrm{BO}}+\hat{H}^{(2)}_{\mathrm{ENC}}-i\hbar\tfrac{d}{dt})|\Phi_{\underline{\underline{\textbf{R}}}}(t)\rangle_{e}$ are the time-dependent vector and scalar potentials, respectively. Here, $\dot{f}$ represents the total time-derivative $df/dt$. The first and second electron–nuclear correlation (ENC) operators are $\hat{H}^{(1)}_{\mathrm{ENC}}=-\sum_{\nu}\dfrac{\hbar^{2}}{2M_{\nu}}\bm{\mathcal{P}}_{\nu}\cdot\nabla_{\nu}\hat{\Gamma}_{\underline{\underline{\textbf{R}}}}$ and $\hat{H}^{(2)}_{\mathrm{ENC}}=-\sum_{\nu}\dfrac{\hbar^{2}}{2M_{\nu}}\left(\nabla_{\nu}^{2}\hat{\Gamma}_{\underline{\underline{\textbf{R}}}}+\nabla_{\nu}\hat{\Gamma}_{\underline{\underline{\textbf{R}}}}\cdot\nabla_{\nu}\hat{\Gamma}_{\underline{\underline{\textbf{R}}}}\right)$ where $\bm{\mathcal{P}}_{\nu}=\nabla_{\nu}|\chi|^{2}/|\chi|^{2}$ is the total nuclear quantum momentum (QM), and $\hat{\Gamma}_{\underline{\underline{\textbf{R}}}}=|\Phi_{\underline{\underline{\textbf{R}}}}\rangle\langle\Phi_{\underline{\underline{\textbf{R}}}}|$ is the reduced electron density operator at a nuclear configuration $\underline{\underline{\textbf{R}}}$}. We note that all terms are evaluated along ${\underline{\underline{\textbf{R}}}}(t)$.
Further expansion of Eq. (2) in terms of BO basis states $|\Phi_{{\underline{\underline{\textbf{R}}}}}(t)\rangle=\sum_{j}C_{j}|\phi_{j}\rangle=\sum_{j}|C_{j}|e^{iS_{j}/\hbar}|\phi_{j}\rangle$ satisfying $\hat{H}_{\mathrm{BO}}|\phi_{j}\rangle=\epsilon_{j}|\phi_{j}\rangle$ yields
@@
For the time evolution of BO coefficients, we have contributions from the Ehrenfest term $\dot{C}_{j}^{\mathrm{Eh}}=-\frac{i}{\hbar}\epsilon_{j}C_{j}-\sum_{\nu,k}\frac{\textbf{P}_{\nu}}{M_{\nu}}\cdot\textbf{d}_{\nu,jk}C_{k}$, the QM term $\dot{C}_{j}^{\mathrm{QM}}=-\sum_{\nu,k}\frac{\bm{\mathcal{P}}_{\nu}}{2M_{\nu}}\cdot\mathbf{D}_{\nu,jk}|C_{k}|^{2}C_{j}$ with $\mathbf{D}_{\nu,jk}=\nabla_{\nu}S_{j}-\nabla_{\nu}S_{k}$, the PQM [32] term $\dot{C}_{j}^{\mathrm{PQM}}=-\sum_{\nu,k}\frac{\bm{\mathcal{Q}}_{\nu,jk}}{2M_{\nu}}\cdot\mathbf{D}_{\nu,jk}|C_{k}|^{2}C_{j}$ where $\bm{\mathcal{Q}}_{\nu,jk}=\nabla_{\nu}|C_{j}|^{2}/|C_{j}|^{2}+\nabla_{\nu}|C_{k}|^{2}/|C_{k}|^{2}$, the divergence term $\dot{C}_{j}^{\mathrm{Div}}=-\sum_{\nu,k}\frac{\nabla_{\nu}\cdot\mathbf{D}_{\nu,jk}}{2M_{\nu}}|C_{k}|^{2}C_{j}$, and the phase correction term $\dot{C}_{j}^{\mathrm{Ph}}=\frac{i}{\hbar}(\tilde{\epsilon}-\sum_{\nu}\frac{|\nabla_{\nu}S_{j}-\textbf{P}_{\nu}|^{2}}{2M_{\nu}})C_{j}$. Here we neglect the $O(\hbar^{2})$ terms and the nonadiabatic coupling (NAC) terms ($\textbf{d}_{\nu,jk}=\langle\phi_{j}|\nabla_{\nu}\phi_{k}\rangle_{e}$ and $g_{\nu,jk}=\langle\phi_{j}|\nabla^{2}_{\nu}\phi_{k}\rangle_{e}$) in ENC terms.
The nuclear force [Eq. (1)] in the BO representation can be expressed as
$\mathbf{F}_{\nu}=\mathbf{F}_{\nu}^{\mathrm{Eh}}+\mathbf{F}_{\nu}^{\mathrm{QM}}+\mathbf{F}_{\nu}^{\mathrm{PQM}}+\mathbf{F}_{\nu}^{\mathrm{Div}}+\mathbf{F}_{\nu}^{\mathrm{Ph}}+\mathbf{F}_{\nu}^{\mathrm{GI}}$ (4)
@@
which directly corresponds to the contributions in Eq. (3) except $\mathbf{F}_{\nu}^{\mathrm{GI}}$ from the gauge-invariant part of the time-dependent scalar potential, $\tilde{\epsilon}_{\mathrm{GI}}=\langle\Phi_{\underline{\underline{\textbf{R}}}}|\hat{H}^{(2)}_{\mathrm{ENC}}|\Phi_{\underline{\underline{\textbf{R}}}}\rangle_{e}$ in Eq. (1). Detailed derivations for Eqs. (3) and (4) are provided in Supplemental Material [33].
@@
The conventional XF-based MQC approaches, coupled-trajectory MQC (CTMQC) [20, 21, 22, 23, 24, 25, 26] or independent-trajectory approaches such as surface hopping dynamics based on XF (SHXF) [27] and Ehrenfest dynamics based on XF (EhXF) [28], only consider $\dot{C}_{j}=\dot{C}^{\mathrm{Eh}}_{j}+\dot{C}^{\mathrm{QM}}_{j}$ and $\mathbf{F}_{\nu}=\mathbf{F}_{\nu}^{\mathrm{Eh}}+\mathbf{F}_{\nu}^{\mathrm{QM}}$ obtained from the first ENC operator $\hat{H}_{\mathrm{ENC}}^{(1)}$ where $\dot{C}^{\mathrm{QM}}_{j}$ and $\mathbf{F}_{\nu}^{\mathrm{QM}}$ provide decoherence with nuclear wave-packet splitting. However, these approaches cannot provide accurate phase evolution [21, 28]. Recently, Arribas and Maitra proposed that an additional correction to the total QM, $\dot{C}^{\mathrm{PQM}}_{j}$, is necessary to consider electronic coherence more accurately [32].
@@
Here we claim that there are more additional terms to be considered at the same or less order as the QM correction, the phase term, $\dot{C}^{\mathrm{Ph}}_{j}$, and the divergence term, $\dot{C}^{\mathrm{Div}}_{j}$ which are crucial to describe correct BO phases and coherence simultaneously. These terms as well as $\dot{C}^{\mathrm{PQM}}_{j}$ are derived from the second ENC operator which has been previously overlooked, especially from $\nabla_{\nu}^{2}\hat{\Gamma}_{\underline{\underline{\textbf{R}}}}$ in $\hat{H}^{(2)}_{\mathrm{ENC}}$ [33].
The phase term, $\dot{C}^{\mathrm{Ph}}_{j}$, is of the same order as $\frac{i}{\hbar}\epsilon_{j}C_{j}$ in $\dot{C}^{\mathrm{Eh}}_{j}$, which governs an additional phase evolution. Here, $\tilde{\epsilon}$ is a global phase, and can therefore be neglected during time evolution. If we approximate $\nabla_{\nu}S_{i}$ by the BO momentum $\textbf{P}_{\nu,i}\equiv\alpha_{i}\textbf{P}_{\nu}$, thus $\mathbf{D}_{\nu,jk}=\textbf{P}_{\nu,j}-\textbf{P}_{\nu,k}$, with $\alpha_{j}$ determined by the state-wise total energy conservation condition, $\alpha_{i}=\sqrt{\max\left(0,\frac{E_{\mathrm{tot}}-\epsilon_{i}}{\sum_{\mu}|\textbf{P}_{\mu,i}|^{2}/M_{\mu}}\right)}$, we obtain
$\begin{split}\dot{S}_{i}-\dot{S}_{j}=&-\sum_{\nu}\dfrac{\textbf{P}_{\nu,j}-\textbf{P}_{\nu,k}}{M_{\nu}}\cdot\textbf{P}_{\nu}\end{split}$ (5)
@@
from phases in $\dot{C}^{\mathrm{Eh}}$ and $\dot{C}^{\mathrm{Ph}}$. This shows that the semiclassical phase correction that had previously been introduced in PCSH heuristically [15] is actually an approximation of the phase term of XF. In this context, we expect that the Stückelberg oscillations can be captured with $\dot{C}^{\mathrm{Ph}}_{j}$ as PCSH.
@@
The divergence contribution, $\dot{C}^{\mathrm{Div}}_{j}$, contains the divergence $\mathbf{D}_{\nu,jk}$, which requires the second derivatives of phases. Despite the difficulty to obtain the second derivatives, the $\dot{C}^{\mathrm{Div}}_{j}$ is essential for the conservation of BO population in the absence of NAC, i.e. $\int d{\underline{\underline{\textbf{R}}}}\dot{|C_{j}|^{2}}|\chi|^{2}=\int\nabla_{\nu}\cdot(\sum_{k}|C_{j}|^{2}|C_{k}|^{2}|\chi|^{2}\mathbf{D}_{\nu,jk})=0$, which is obtained from $\dot{C}^{\mathrm{QM}}+\dot{C}^{\mathrm{PQM}}+\dot{C}^{\mathrm{Div}}$ since $\dot{C}^{\mathrm{Eh}}$ without NAC and $\dot{C}^{\mathrm{Ph}}$ does not contribute to the BO population transfer. Therefore, in this work, we approximate the divergence term using constant pair-wise shift terms (See Supplemental Material [33] for the numerical implementation of the equations).
@@
We benchmark the new equations of motion using the one-dimensional DAG [14] and the 2DNS [15] models. The model Hamiltonians and computational details are provided in the Supplemental Material [33].
@@
Fig. 1 shows the transmission probability on the lower state, $T_{1}$, for the DAG model with various trajectory-based approaches in comparison with the exact wave-packet dynamics. We refer to the results obtained without the additional terms as SHXF and CT, whereas those including the term are denoted as SHXFv2 and CTv2, respectively. The exact behavior shows an increase of oscillation amplitude as $k_{0}$ increases, from $\sim$ 0.5 at $k_{0}=30.0$ a.u. As reported previously, PCSH can correctly reproduce the exact Stückelberg oscillation as aimed [15]. On the other hand, PCSH with the energy-based decoherence correction [11] (PCSH-EDC) removes the oscillation due to over-decoherence. Similarly, CT, which only account for the decoherence, shows $T_{1}\sim 0.5$ for all values of $k_{0}$. Although SHXF shows an oscillatory behavior with respect to $k_{0}$, it deviates from the exact behavior. CTv2 and SHXFv2 recover the correct Stückelberg oscillation. While CTv2 slightly underestimates the overall amplitude, SHXFv2 shows excellent agreement with the exact result at high $k_{0}$, and is qualitatively correct at low $k_{0}$. This demonstrates that the additional phase correction term is essential for the correct time evolution of BO phases.
Figure 2: Time evolution of BO population $\langle|C_{1}|^{2}\rangle(t)$ (upper) and coherence $\langle|C_{1}C_{2}|^{2}\rangle(t)$ (lower) with different initial momenta $k_{0}=30$ and $40$ a.u.
Although reproducing the correct oscillation in $T_{1}$ is encouraging, it remains essential to evaluate the time propagation of wave packets to fully assess the accuracy. Fig. 2 shows the time evolution of BO population for the lower state, $\langle|C_{1}|^{2}\rangle(t)$, and coherence, $\langle|C_{1}C_{2}|^{2}\rangle(t)$, with different initial momenta $k_{0}=30.0$ and $50.0$ a.u. For exact wave-packet dynamics, $\langle O\rangle(t)$ is defined as $\langle O\rangle(t)=\int d{\underline{\underline{\textbf{R}}}}\;O({\underline{\underline{\textbf{R}}}},t)$, whereas $\langle O\rangle(t)\equiv\sum_{I}^{N_{\mathrm{traj}}}O^{(I)}(t)/N_{\mathrm{traj}}$ is used for MQC methods. Since the DAG model contains two NAC regions, the initial population drop corresponds to passage through the first region, while the separated wave packets on upper and lower states afterward encounter the second region at different times.

### v2

@@ Abstract
We propose mixed quantum–classical equations of motion that unify electronic coherence and phase evolution simultaneously within the exact factorization framework. Our derivation shows that incorporating the second-order electron–nuclear correlation terms from the exact coupled time-dependent Schrödinger equations is essential to recover both correct phase dynamics and complete electronic (de)coherence, including their effect on nuclear forces. Benchmark calculations on one- and two-dimensional model systems confirm that the approach accurately captures key nonadiabatic features. The equations therefore provide a rigorous first-principles foundation for mixed quantum–classical description of coupled electron–nuclear dynamics, bringing electronic coherence and phase evolution—long treated through separate heuristic corrections—into a single unified and systematically derived framework.
@@
Understanding the coupled motion of electrons and nuclei in molecules remains one of the central challenges in theoretical chemistry and chemical physics. Fully quantum treatments of electron–nuclear dynamics are computationally prohibitive for systems of realistic size, owing to the scaling of the many-body molecular wavefunction. To overcome this limitation, mixed quantum–classical (MQC) methods [1, 2, 3] where the electrons are treated quantum mechanically while nuclei follow classical trajectories have become indispensable tools for simulating photochemical processes, charge and energy transfer, and ultrafast dynamics in molecules and materials [4, 5, 6, 7, 8, 9]. Among MQC methods, Ehrenfest dynamics [1] and trajectory-based surface hopping [2] are the most widely applied. Despite their utility, they suffer from two fundamental shortcomings: the absence of a reliable mechanism for electronic decoherence, and the lack of a rigorous treatment of phase evolution between adiabatic states. Numerous correction schemes for decoherence and phase evolution have been proposed, often guided by semiclassical arguments, empirical adjustments, intuition from wave-packet models, or mapping of variables [10, 11, 12, 13, 14, 15, 16, 17]. Many approaches have successfully described electronic coherence or phase evolution, but they usually address only one of them or overlook their time-resolved nuclear wave-packet and electronic dynamics.
@@
The exact factorization (XF) formalism provides a rigorous foundation for electron–nuclear dynamics by expressing the molecular wavefunction as a product of a nuclear wavefunction and a conditional electronic state [18, 19]. This leads to coupled electronic and nuclear equations of motion that, in principle, include the full electron–nuclear correlation (ENC) effects. XF-based approaches have therefore attracted attention as a systematic framework for deriving improved MQC dynamics [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. In recent work, Arribas and Maitra showed that a projected quantum momentum (PQM) correction naturally arises from the XF formalism and plays a central role in describing electronic coherence [33]. However, the correct phase-evolution effect in MQC dynamics has not been identified from the electron–nuclear correlation term within the XF framework. As a consequence, XF-based MQC methods fail to reproduce the proper relative phase evolution of the adiabatic states and, in turn, an important nonadiabatic feature, namely the Stückelberg oscillation.
@@
In this work, we derive new mixed quantum–classical equations of motion from the exact factorization formalism by including the terms from the previously neglected second-order ENC operator, revealing an additional phase-correction term that accompanies the PQM correction of comparable or smaller order in $\hbar$. This previously unidentified term governs the proper phase evolution of the Born–Oppenheimer (BO) coefficients. The combined inclusion of PQM and phase corrections provides a balanced and rigorous description of both decoherence and phase dynamics without relying on semiclassical approximations or heuristic arguments, offering deeper conceptual insight into ENC and a practical route to accurate molecular quantum dynamics. We validate the equations through benchmark calculations for the double-arch geometry (DAG) [14], dual avoided crossing (DAC) [2], and two-dimensional nonseparable (2DNS) [15] models, which demonstrate that both corrections are essential for reproducing characteristic nonadiabatic features such as Stückelberg oscillations, intermediate-time electronic coherence, and the correct nuclear-density distributions.
@@
where $\hat{H}_{\mathrm{BO}}$ is the BO Hamiltonian, and $\textbf{A}_{\nu}({\underline{\underline{\textbf{R}}}},t)=\langle\Phi_{\underline{\underline{\textbf{R}}}}|-i\hbar\nabla_{\nu}\Phi_{\underline{\underline{\textbf{R}}}}\rangle_{e}$ and $\tilde{\epsilon}({\underline{\underline{\textbf{R}}}},t)=\langle\Phi_{\underline{\underline{\textbf{R}}}}(t)|(\hat{H}_{\mathrm{BO}}+\hat{H}^{(2)}_{\mathrm{ENC}}-i\hbar\tfrac{d}{dt})|\Phi_{\underline{\underline{\textbf{R}}}}(t)\rangle_{e}$ are the time-dependent vector and scalar potentials, respectively. Here, $\dot{f}$ represents $df/dt=\partial f/\partial t+\sum_{\nu}\dot{\textbf{R}}_{\nu}\cdot\nabla_{\nu}f$. The first- and second-order electron–nuclear correlation (ENC) operators are
$\hat{H}^{(1)}_{\mathrm{ENC}}=-\sum_{\nu}\dfrac{\hbar^{2}}{2M_{\nu}}\bm{\mathcal{P}}_{\nu}\cdot\nabla_{\nu}\hat{\Gamma}_{\underline{\underline{\textbf{R}}}}$ (3)
and
$\hat{H}^{(2)}_{\mathrm{ENC}}=-\sum_{\nu}\dfrac{\hbar^{2}}{2M_{\nu}}\left(\nabla_{\nu}^{2}\hat{\Gamma}_{\underline{\underline{\textbf{R}}}}+\nabla_{\nu}\hat{\Gamma}_{\underline{\underline{\textbf{R}}}}\cdot\nabla_{\nu}\hat{\Gamma}_{\underline{\underline{\textbf{R}}}}\right),$ (4)
respectively, where $\bm{\mathcal{P}}_{\nu}=\nabla_{\nu}|\chi|^{2}/|\chi|^{2}$ is the total nuclear quantum momentum (QM), and $\hat{\Gamma}_{\underline{\underline{\textbf{R}}}}=|\Phi_{\underline{\underline{\textbf{R}}}}\rangle\langle\Phi_{\underline{\underline{\textbf{R}}}}|$ is the reduced electron density operator at a nuclear configuration $\underline{\underline{\textbf{R}}}$}. We note that all terms are evaluated along ${\underline{\underline{\textbf{R}}}}(t)$ and the above equations are written in a general gauge without any specific choice of gauge.
Further expansion of eq 2 in terms of BO basis states $|\Phi_{{\underline{\underline{\textbf{R}}}}}(t)\rangle=\sum_{j}C_{j}|\phi_{j}\rangle=\sum_{j}|C_{j}|e^{iS_{j}/\hbar}|\phi_{j}\rangle$ satisfying $\hat{H}_{\mathrm{BO}}|\phi_{j}\rangle=\epsilon_{j}|\phi_{j}\rangle$ yields
@@
For the time evolution of BO coefficients, we have contributions from the Ehrenfest term $\dot{C}_{j}^{\mathrm{Eh}}=-\frac{i}{\hbar}\epsilon_{j}C_{j}-\sum_{\nu,k}\frac{\textbf{P}_{\nu}}{M_{\nu}}\cdot\textbf{d}_{\nu,jk}C_{k}$, the QM term $\dot{C}_{j}^{\mathrm{QM}}=-\sum_{\nu,k}\frac{\bm{\mathcal{P}}_{\nu}}{2M_{\nu}}\cdot\mathbf{D}_{\nu,jk}|C_{k}|^{2}C_{j}$ with $\mathbf{D}_{\nu,jk}=\nabla_{\nu}S_{j}-\nabla_{\nu}S_{k}$, the PQM [33] term
$\dot{C}_{j}^{\mathrm{PQM}}=-\sum_{\nu,k}\frac{\bm{\mathcal{Q}}_{\nu,jk}}{2M_{\nu}}\cdot\mathbf{D}_{\nu,jk}|C_{k}|^{2}C_{j}$ (6)
where $\bm{\mathcal{Q}}_{\nu,jk}=\nabla_{\nu}|C_{j}|^{2}/|C_{j}|^{2}+\nabla_{\nu}|C_{k}|^{2}/|C_{k}|^{2}$, the divergence term
$\dot{C}_{j}^{\mathrm{Div}}=-\sum_{\nu,k}\frac{\nabla_{\nu}\cdot\mathbf{D}_{\nu,jk}}{2M_{\nu}}|C_{k}|^{2}C_{j},$ (7)
and the phase correction term
$\dot{C}_{j}^{\mathrm{Ph}}=\frac{i}{\hbar}(\tilde{\epsilon}-\sum_{\nu}\frac{|\nabla_{\nu}S_{j}-\textbf{P}_{\nu}|^{2}}{2M_{\nu}})C_{j}.$ (8)
Here we neglect the $O(\hbar^{2})$ terms and the nonadiabatic coupling (NAC) terms ($\textbf{d}_{\nu,jk}=\langle\phi_{j}|\nabla_{\nu}\phi_{k}\rangle_{e}$ and $g_{\nu,jk}=\langle\phi_{j}|\nabla^{2}_{\nu}\phi_{k}\rangle_{e}$) in ENC terms.
The nuclear force (eq 1) in the BO representation can be expressed as
$\mathbf{F}_{\nu}=\mathbf{F}_{\nu}^{\mathrm{Eh}}+\mathbf{F}_{\nu}^{\mathrm{QM}}+\mathbf{F}_{\nu}^{\mathrm{PQM}}+\mathbf{F}_{\nu}^{\mathrm{Div}}+\mathbf{F}_{\nu}^{\mathrm{Ph}}+\mathbf{F}_{\nu}^{\mathrm{GI}},$ (9)
where $\mathbf{F}_{\nu}^{\mathrm{Eh}}=-\sum_{k}|C_{k}|^{2}\nabla_{\nu}\epsilon_{k}-\sum_{k,l}C_{k}C_{l}^{*}(\epsilon_{k}-\epsilon_{l})\textbf{d}_{\nu,lk}$, $\mathbf{F}_{\nu}^{\mathrm{QM}}=-\sum_{k,l}\mathbf{D}_{\nu,kl}\left(\sum_{\mu}\frac{\bm{\mathcal{P}}_{\mu}}{2M_{\mu}}\cdot\mathbf{D}_{\mu,kl}\right)|C_{l}|^{2}|C_{k}|^{2}$,
$\mathbf{F}_{\nu}^{\mathrm{PQM}}=-\sum_{k,l}\mathbf{D}_{\nu,kl}\left(\sum_{\mu}\dfrac{\bm{\mathcal{Q}}_{\mu,kl}}{2M_{\mu}}\cdot\mathbf{D}_{\mu,kl}\right)|C_{l}|^{2}|C_{k}|^{2},$ (10)
$\mathbf{F}_{\nu}^{\mathrm{Div}}=-\sum_{k,l}\mathbf{D}_{\nu,kl}\left(\sum_{\mu}\dfrac{\nabla_{\mu}}{2M_{\mu}}\cdot\mathbf{D}_{\mu,kl}\right)|C_{l}|^{2}|C_{k}|^{2},$ (11)
and
$\mathbf{F}_{\nu}^{\mathrm{Ph}}+\mathbf{F}_{\nu}^{\mathrm{GI}}=-\sum_{k,l}\sum_{\mu}\left(\dfrac{\mathbf{H}_{\nu\mu,kl}\mathbf{D}_{\mu,kl}}{2M_{\mu}}\right)|C_{l}|^{2}|C_{k}|^{2},$ (12)
@@
where $\mathbf{H}_{\nu\mu,kl}$ is a $3\times 3$ cartesian sub-Hessian matrix of phase difference for nuclei $\nu$ and $\mu$, i.e. $\left[\mathbf{H}_{\nu\mu,kl}\right]_{pq}=\partial^{2}/\partial p_{\nu}\partial q_{\mu}(S_{k}-S_{l})$ for cartesian components $p$ and $q$. Each term directly corresponds to the contributions in eq 5 except $\mathbf{F}_{\nu}^{\mathrm{GI}}$ from the gauge-invariant part of the time-dependent scalar potential, $\tilde{\epsilon}_{\mathrm{GI}}=\langle\Phi_{\underline{\underline{\textbf{R}}}}|\hat{H}^{(2)}_{\mathrm{ENC}}|\Phi_{\underline{\underline{\textbf{R}}}}\rangle_{e}$ in eq 1, where a part of $\mathbf{F}_{\nu}^{\mathrm{GI}}$ cancels with $\mathbf{F}_{\nu}^{\mathrm{Ph}}$ to generate the net force in eq 12. Detailed derivations are provided in Supporting Information.
@@
The conventional XF-based MQC approaches, coupled-trajectory MQC (CTMQC) [20, 21, 22, 23, 24, 25, 26] or independent-trajectory approaches such as surface hopping dynamics based on XF (SHXF) [27] and Ehrenfest dynamics based on XF (EhXF) [28], only consider $\dot{C}_{j}=\dot{C}^{\mathrm{Eh}}_{j}+\dot{C}^{\mathrm{QM}}_{j}$ and $\mathbf{F}_{\nu}=\mathbf{F}_{\nu}^{\mathrm{Eh}}+\mathbf{F}_{\nu}^{\mathrm{QM}}$ obtained from the first-order ENC operator $\hat{H}_{\mathrm{ENC}}^{(1)}$ where $\dot{C}^{\mathrm{QM}}_{j}$ and $\mathbf{F}_{\nu}^{\mathrm{QM}}$ provide decoherence with nuclear wave-packet splitting. However, these approaches cannot provide accurate phase evolution [21, 28]. Recently, Arribas and Maitra proposed that an additional correction to the total QM, $\dot{C}^{\mathrm{PQM}}_{j}$, is necessary to consider electronic coherence more accurately [33].
@@
Here we claim that there are more additional terms to be considered at the same or less order in $\hbar$ as the QM correction, the phase term, $\dot{C}^{\mathrm{Ph}}_{j}$, and the divergence term, $\dot{C}^{\mathrm{Div}}_{j}$ which are crucial to describe correct BO phases and coherence simultaneously. These terms as well as $\dot{C}^{\mathrm{PQM}}_{j}$ are derived from the second ENC operator which has been previously overlooked, especially from the $\nabla_{\nu}^{2}\hat{\Gamma}_{\underline{\underline{\textbf{R}}}}$ term in $\hat{H}^{(2)}_{\mathrm{ENC}}$.
[…]
$\begin{split}\dot{S}_{j}-\dot{S}_{k}=&-\sum_{\nu}\dfrac{\textbf{P}_{\nu,j}-\textbf{P}_{\nu,k}}{M_{\nu}}\cdot\textbf{P}_{\nu}\end{split}$ (13)
@@
from phases in $\dot{C}_{j}^{\mathrm{Eh}}$ and $\dot{C}_{j}^{\mathrm{Ph}}$. This shows that the semiclassical phase correction that had previously been introduced in phase-corrected surface hopping (PCSH) heuristically [15] is actually an approximation of the phase term of XF. In this context, we expect that the Stückelberg oscillations can be captured with $\dot{C}^{\mathrm{Ph}}_{j}$ as they are in PCSH.
@@
The divergence contribution, $\dot{C}^{\mathrm{Div}}_{j}$, contains the divergence of $\mathbf{D}_{\nu,jk}$, which requires the second derivatives of phases. Despite the difficulty to obtain the second derivatives, the $\dot{C}^{\mathrm{Div}}_{j}$ is essential for the conservation of BO population in the absence of NAC, i.e. $\int d{\underline{\underline{\textbf{R}}}}\dot{|C_{j}|^{2}}|\chi|^{2}=\int\nabla_{\nu}\cdot(\sum_{k}|C_{j}|^{2}|C_{k}|^{2}|\chi|^{2}\mathbf{D}_{\nu,jk})=0$, which is obtained from $\dot{C}_{j}^{\mathrm{QM}}+\dot{C}_{j}^{\mathrm{PQM}}+\dot{C}_{j}^{\mathrm{Div}}$ since $\dot{C}_{j}^{\mathrm{Eh}}$ without NAC and $\dot{C}_{j}^{\mathrm{Ph}}$ does not contribute to the BO population transfer. Therefore, in this work, we approximate the divergence term using constant pair-wise shift terms (See Supporting Information for the numerical implementation of the equations).
@@
We benchmark the new equations of motion using the one-dimensional DAG [14], DAC [2] and the 2DNS [15] models. The model Hamiltonians and computational details are provided in the Supporting Information.
@@
Figure 1 shows the transmission probability on the lower state, $T_{1}$, for the DAG model with various trajectory-based approaches in comparison with the exact wave-packet dynamics. We refer to the results obtained without the additional terms as SHXF and CT, whereas those including the terms are denoted as SHXFv2 and CTv2, respectively. The exact behavior shows an increase of oscillation amplitude as $k_{0}$ increases, from $\sim$ 0.5 at $k_{0}=30.0$ a.u. As reported previously, PCSH can correctly reproduce the exact Stückelberg oscillation as aimed [15]. On the other hand, PCSH with the energy-based decoherence correction [11] (PCSH-EDC) removes the oscillation due to over-decoherence. Similarly, CT, which only account for the decoherence, shows $T_{1}\sim 0.5$ for all values of $k_{0}$. Although SHXF shows an oscillatory behavior with respect to $k_{0}$, it deviates from the exact behavior. CTv2 and SHXFv2 recover the correct Stückelberg oscillation. While CTv2 slightly underestimates the overall amplitude, SHXFv2 shows excellent agreement with the exact result at high $k_{0}$, and is qualitatively correct at low $k_{0}$. This demonstrates that the additional phase correction term is essential for the correct time evolution of BO phases. Based on the results in Figures LABEL:fig:pop_coh_35LABEL:fig:ct2_spatial_k0_35, and LABEL:fig:shxf2_spatial_k0_35, we attribute the underestimation of the oscillation amplitude in the low-$k_{0}$ region to the numerical approximations rather than to the over-decoherence, although this needs to be further investigated in detail for more accurate simulations in the future. Similarly, the DAC model, a standard Tully model which suffers from phase problems, shows that CTv2 and SHXFv2 well reproduce the correct T1 (Figure LABEL:fig:RT_dac).
Figure 2: Time evolution of BO population $\langle|C_{1}|^{2}\rangle(t)$ (upper) and coherence $\langle|C_{1}C_{2}|^{2}\rangle(t)$ (lower) with different initial momenta $k_{0}=30.0$ and $40.0$ a.u.
Although reproducing the correct oscillation in $T_{1}$ is encouraging, it remains essential to evaluate the time propagation of wave packets to fully assess the accuracy. Figure 2 shows the time evolution of BO population for the lower state, $\langle|C_{1}|^{2}\rangle(t)$, and coherence, $\langle|C_{1}C_{2}|^{2}\rangle(t)$, with different initial momenta $k_{0}=30.0$ and 40.0 a.u. For exact wave-packet dynamics, $\langle O\rangle(t)$ is defined as $\langle O\rangle(t)=\int d{\underline{\underline{\textbf{R}}}}\;O({\underline{\underline{\textbf{R}}}},t)|\chi({\underline{\underline{\textbf{R}}}},t)|^{2}$, whereas $\langle O\rangle(t)\equiv\sum_{I}^{N_{\mathrm{traj}}}O^{(I)}(t)/N_{\mathrm{traj}}$ is used for MQC methods. Since the DAG model contains two NAC regions, the initial population drop corresponds to passage through the first region, while the separated wave packets on upper and lower states afterward encounter the second region at different times.